# Augmenting Human Expertise in Weighted Ensemble Simulations through Deep Learning based Information Bottleneck

### v1

@@ I Introduction
Advances in computational hardware and algorithms have made all-atom molecular dynamics (MD) a key tool for studying chemical and biological systems. By implementing physical laws in silico, MD simulations allow for detailed tracking of molecular evolution over large spatiotemporal domains—up to thousands of angstroms with atomic precision and up to milliseconds at femtosecond resolution.Dror et al. (2012); Hollingsworth and Dror (2018) However, MD simulations have their challenges. Many important processes, such as ligand dissociation or protein conformation changes, occur over timescales ranging from milliseconds to hundreds of seconds and involve very high free energy barriers. MD simulations often struggle to overcome these barriers, making such events exceedingly rare and difficult to capture. Even with the most advanced resources available, simulating these processes can still take months or even years.Shaw et al. (2021)
@@
A variety of methods, generally called enhanced sampling techniques, have emerged as a solution to the rare event problem. For a comprehensive review of the enhanced sampling methods used in molecular dynamics simulations, interested readers can refer to Refs. Bernardi, Melo, and Schulten, 2015; Yang et al., 2019; Hénin et al., 2022; Mehdi et al., 2024. In this work, we majorly focus on an increasingly popular segment-based sampling method called weighted ensemble (WE).Huber and Kim (1996); Zuckerman and Chong (2017) The WE method has demonstrated its versatility across a spectrum of challenges, from protein folding,Adhikari et al. (2019) protein-protein binding,Saglam and Chong (2019) and protein-ligand unbinding,Lotz and Dickson (2018) to the large-scale dynamics of the SARS-CoV-2 spike protein.Sztain et al. (2021) Motivated by these real-world applications, the field is undergoing rapid evolution, which is further catalyzed by the emergence of easy-to-use and developer-friendly tools such as WESTPA 2.0.Russo et al. (2022)
@@
At the heart of the WE framework lies the process of mapping the configuration space onto a low-dimensional collective variable (CV) space, followed by partitioning it into bins. This mapping and binning process enables the dynamic splitting and merging of trajectories based on their progress along the identified CVs, thereby enhancing the sampling of rare events. The selection of appropriate CVs and binning schemes can significantly influence the algorithm’s performance. Voronoi polyhedra offer a convenient method to tessellate spaces of any dimensionality. This technique has been leveraged to enable adaptive binning in a hierarchical manner for the WE method.Zhang, Jasnow, and Zuckerman (2010); Dickson and Brooks III (2014) Based on the finite-temperature string method, Ref. Adelman and Grabe, 2013 adaptively learns a string of Voronoi cells to restrict the WE sampling to a one-dimensional path embedded in a multi-dimensional space defined by some CVs. Later, a minimal adaptive binning (MAB) schemeTorrillo, Bogetti, and Chong (2021) is proposed to automate the placement of bins along a chosen CV, which effectively surmounts large free energy barriers encountered during the WE simulation. Alternatively, recent advancements propose optimizing both bin construction and trajectory allocation based on a newly developed mathematical formulation.Aristoff et al. (2023) Despite these strides towards more automatic binning schemes, the determination of CVs in WE simulations still heavily relies on expert intuition.
@@
Moreover, as an unbiased method, the WE method suffers from slow convergence. To speed up the convergence, Ref. Ahn et al., 2021 uses Gaussian accelerated molecular dynamics to generate an initial set of configurations. These configurations, weighted appropriately, are then utilized to seed the WE simulations, enhancing convergence. Similarly, Ref. Ojha et al., 2023 adopts a comparable approach, enhancing the initial state distribution for WE simulations by constructing a Markov state model (MSM) from short unbiased trajectories. Moreover, history-augmented Markov state models (haMSMs) have emerged as potent tools for estimating stationary distributions and rate constants from transient, unconverged WE data.Copperman and Zuckerman (2020) These estimates can then also serve as effective starting points for initiating new WE simulations, thereby accelerating the convergence of the simulation.
@@ II.1 Weighted Ensemble (WE) Algorithm
In a typical MD simulation, the system spends most of its time in high-probability basins, and rarely leaves the basins to visit low-probability regions such as transition and metastable states. The WE algorithm Huber and Kim (1996); Zuckerman and Chong (2017) adopts a splitting strategy using an ensemble of system copies (or “replicas”), each of which carries a statistical weight. These replicas are split or merged at a periodic time interval, called the resampling time $\tau$, guaranteeing an even coverage of the configuration space. Through this resampling procedure, the WE method is capable of enhancing the sampling of the low-probability regions. As these replicas are assigned statistical weights in a rigorous manner, in principle no bias is introduced in the dynamics. Thus, the WE algorithm can yield unbiased estimates of both the thermodynamic and kinetic properties of the molecular system.
@@
Figure 1: An illustration of the hybrid approach combining expert-based CVs and data-driven SPIB-learned CVs for two instances of a protein conformation (top and bottom). Replicas are first mapped both to the expert-based and SPIB-learned CVs, denoted with subscripts $e$ and $s$ respectively. The top (bottom) replica is mapped to bin $i_{e}$ ($i^{\prime}_{e}$) based on expert-based bins and bin $i_{s}$ ($i^{\prime}_{s}$) based on SPIB-learned bins, as indicated by the arrows. Undersampled regions, defined by expert-based bins, are shaded in the bottom right figure. The total number of replicas mapped to these bins is represented by the number of circles. Since bin $i_{e}$ is not in an undersampled region, the top structure is assigned to bin $i_{s}$ (solid magenta arrow), regardless of the number of replicas in bins $i_{e}$ and $i_{s}$. However, since bin $i^{\prime}_{e}$ is in an undersampled region and contains fewer replicas (with only one replica) compared to bin $i^{\prime}_{s}$ (which has two replicas), the bottom replica is assigned to $i^{\prime}_{e}$ (solid green arrow).
@@ II.2 State Predictive Information Bottleneck (SPIB)
State predictive information bottleneck (SPIB)Wang and Tiwary (2021); Wang et al. (2024) uses an information bottleneck-based protocol to learn CVs of desired dimensionality from time resolved trajectory data, while simultaneously identifying important metastable states. Previous studies have demonstrated its efficacy in analyzing complex molecular systems by constructing multi-resolution MSMs.Wang et al. (2022, 2024) Furthermore, SPIB has proven valuable in learning low-dimensional CVs for enhanced sampling methods. This approach accelerates diverse processes, including ligand permeationMehdi et al. (2022) and dissociation,Beyerle and Tiwary ; Lee et al. (2024) protein conformational changes,Vani et al. (2023); Vani, Aranganathan, and Tiwary (2023); Gu, Aranganathan, and Tiwary (2024) and crystal polymorph nucleation.Zou et al. (2023); Wang et al. (2023)
@@
SPIB adopts a structure akin to a variational autoencoder (VAE),Kingma and Welling (2013) consisting of two neural networks: An encoder maps the input features, such as those characterizing a protein’s current configuration, to a low-dimensional CV space, while a decoder tries to predict some target from this space. However, unlike VAE, where the aim is to reconstruct the entire input, SPIB only require the learned CVs to predict coarse-grained future state labels, namely which state the system will occupy after a pre-specified lag time $\Delta t$. The main advantage of such a simplification of the prediction task is that only the motion related to the transitions between different states will be captured by the learned CVs, while the fluctuations inside any metastable state will be ignored. Typically, the number and location of such states are not available a priori, but the SPIB method makes it possible to robustly estimate these on-the-fly.Wang and Tiwary (2021); Wang et al. (2024) In this way, SPIB can autonomously cluster similar MD conformations into metastable states and accurately identify transition pathways between these states. This process allows SPIB to learn a low-dimensional CV space, where distinct states are separated into basins and connected by key transition pathways. Such clusterability of SPIB empowers the learned CVs to prioritize the representation of state-to-state transitions.
@@
By introducing a lag time $\Delta t$, SPIB acts as a “fast mode filter”, selectively identifying metastable states with significant Boltzmann weights and lifetimes surpassing $\Delta t$, while excluding modes equilibrating faster.Wang et al. (2024) This enables SPIB to automatically pinpoint important metastable states in the system. $\Delta t$ serves as the minimum time resolution of interest. Aligning $\Delta t$ with the WE resampling time $\tau$—such as selecting $\Delta t$ as half of $\tau$—allows SPIB to filter out fast processes while retaining slower ones that cannot be sampled within $\tau$, making SPIB ideal for integration into WE simulations.
@@
The CVs learned by SPIB can be interpreted as a low-dimensional continuous embedding of MD conformations that preserves maximal information about state-to-state transitions, effectively capturing all transitions between identified states.Wang et al. (2024) This is unlike commonly used dynamics-based dimension reduction methods such as tICA,Noé and Nuske (2013); Pérez-Hernández et al. (2013) which focus on learning a given number of slow processes. For instance, a 2D tICA captures only the two slowest processes while a 2D SPIB can accurately capture not just the two slowest, but all slow transition processes between the identified metastable states. This inherent distinction renders SPIB suitable for learning a low-dimensional CV space for systems featuring multiple metastable states of interest.
@@ II.3 SPIB Augmented Enhanced Sampling
A solution to this conundrum could involve an iterative approach that alternates between rounds of enhanced sampling and SPIB to gradually improve sampling.Wang et al. (2019); Mehdi et al. (2024) In the Reweighted Autoencoded Variational Bayes (RAVE) method,Ribeiro et al. (2018); Wang, Ribeiro, and Tiwary (2019); Mehdi et al. (2022) the process starts with initial trajectories, and SPIB learns CVs, which may initially be suboptimal, to guide the subsequent enhanced sampling simulation to explore an increased amount of configuration space. Once a better sampling is collected, better CVs can then be learned. This looping of enhanced sampling and SPIB can continue until a satisfactory level of sampling is achieved.
However, as a deep neural network-based method, SPIB usually fails in extrapolation. This means that solely relying on SPIB to guide sampling may lead the system to get trapped in some metastable states, hindering effective exploration of new configurations. One solution to strengthen the extrapolability of such machine learned CVs is to use a linear encoder paired with a non-linear decoder in SPIB, as we have done in many previous applications.Beyerle and Tiwary ; Lee, Wang, and Tiwary (2024); Vani et al. (2023); Vani, Aranganathan, and Tiwary (2023); Gu, Aranganathan, and Tiwary (2024); Zou et al. (2023); Wang et al. (2023)
Here, we aim to explore an alternative approach by directly combining SPIB learned CVs with expert intuition based CVs. This is primarily motivated by recognizing that while experts may not grasp all the intricate dynamics of complex molecular systems, they possess intuition about crucial, more global demarcators of state-to-state transitions. Hence, expert-based CVs like Root-mean-square deviation of atomic positions (RMSD), radius of gyration (Rg), or specific distances and dihedral angles, remain prevalent in many current enhanced sampling studies, despite the emergence of many ML-based CV identification approaches. These expert-based CVs excel in discerning states of interest, exhibit good extrapolation, and provide valuable insights into system dynamics. However, they may not capture all transition details accurately, thus hindering the identification of exact transition pathways for complex systems. Moreover, due to the prohibitive cost associated with sampling high-dimensional spaces, these expert-based CVs also are constrained by their dimensionality, posing additional challenges in proposing effective CVs.
In contrast, ML-based CVs, such as those learned by SPIB, excel in distinguishing states of interest, are inherently low-dimensional, and are adept at capturing dynamic details and identifying transition pathways. Nonetheless, they commonly struggle with extrapolation. Additionally, these ML-based methods are typically versatile, applicable to a wide range of problems and systems. Yet, there’s often a desire to incorporate system-specific intuition to guide exploration and CV identification processes in practical applications.
@@ II.4 A Hybrid Approach for CV construction for Weighted Ensemble
In this work, we adopt a uniform binning scheme for both expert-based CVs and SPIB-learned CVs. Therefore, combining expert-based and SPIB-learned CVs involves merging bins derived from both sets of CVs. A simple strategy is proposed in Fig. 1. For instance, if a replica is assigned to bin $i_{e}$ based on the expert-based bins and to bin $i_{s}$ based on the SPIB-learned bins (where $i$ represents the bin index, and subscripts $e$ and $s$ denote expert-based and SPIB-learned, respectively), we need to determine when to use the expert-based binning and assign the replica to bin $i_{e}$ and when to use the SPIB-learned binning and assign the replica to bin $i_{s}$. First, we define the undersampled regions based on expert-based bins. Utilizing the trajectory data employed for SPIB training, we count the number of samples within each bin by mapping all samples to the expert-based bins, without considering sample weights. Subsequently, expert-based bins with counts below the average are identified as undersampled regions. This determination is updated after each SPIB training. Then, we assess whether the replica is in undersampled regions. If a replica is in undersampled regions and there are fewer replicas mapped to bin $i_{e}$ than to bin $i_{s}$, the replica is assigned to bin $i_{e}$; otherwise, it is assigned to bin $i_{s}$. This hybrid strategy effectively combines the strengths of both SPIB-based CVs and expert-based CVs. In well-sampled regions, SPIB-learned CVs effectively separate metastable states and highlight key transition pathways, offering reliable guidance for enhanced sampling and facilitating slow state-to-state transitions. Conversely, in undersampled regions, the hybrid approach utilizes both SPIB-based bins and expert-based bins to improve sampling of undersampled transition states and promote exploration of new regions. This approach facilitates effective navigation of complex systems, ensuring efficient and robust sampling across the configuration space of interest.
@@
The full workflow is summarized in Fig. 2. Initially, utilizing expert-based CVs and their corresponding binning scheme, $N_{0}$ iterations of WE simulations are conducted to gather initial trajectories. Next, we train the SPIB model on the gathered MD data, clustering the sampled MD conformations into metastable states. This process allows SPIB to effectively map explored regions to a low-dimensional CV space, where distinct states are separated into basins connected by key transition pathways. A more detailed discussion of the neural network architecture and model training is provided in the Supporting Information (SI). In this work, we employ a rectilinear grid with equal bin sizes to uniformly bin the SPIB-learned CV space. We implement a straightforward algorithm to automatically adjust the bin size, aiming to maintain a roughly constant desired number of occupied SPIB bins (see SI for details). These SPIB-based bins, in combination with the expert-based bins, are then used to guide further WE simulations. $N_{\text{update}}$ iterations of WE simulations are conducted using these hybrid bins. Subsequently, based on the newly collected data, a new SPIB model can be obtained. This iterative process between WE simulations and SPIB training continues until a satisfactory level of sampling is achieved.

### v2

@@ I Introduction
Advances in computational hardware and algorithms have made all-atom molecular dynamics (MD) a key tool for studying chemical and biological systems. By implementing physical laws in silico, MD simulations allow for detailed tracking of molecular evolution over large spatiotemporal domains—up to thousands of angstroms with atomic precision and up to milliseconds at femtosecond resolution.[1, 2] However, MD simulations have their challenges. Many important processes, such as ligand dissociation or protein conformation changes, occur over timescales ranging from milliseconds to hundreds of seconds and involve very high free energy barriers. MD simulations often struggle to overcome these barriers, making such events exceedingly rare and difficult to capture. Even with the most advanced resources available, simulating these processes can still take months or even years.[3]
@@
A variety of methods, generally called enhanced sampling techniques, have emerged as a solution to the rare event problem. For a comprehensive review of the enhanced sampling methods used in molecular dynamics simulations, interested readers can refer to Refs. 4, 5, 6, 7. In this work, we majorly focus on an increasingly popular segment-based sampling method called weighted ensemble (WE).[8, 9] The WE method has demonstrated its versatility across a spectrum of challenges, from protein folding,[10] protein-protein binding,[11] and protein-ligand unbinding,[12] to the large-scale dynamics of the SARS-CoV-2 spike protein.[13] Motivated by these real-world applications, the field is undergoing rapid evolution, which is further catalyzed by the emergence of easy-to-use and developer-friendly tools such as WESTPA 2.0.[14]
@@
At the heart of the WE framework lies the process of mapping the configuration space onto a low-dimensional collective variable (CV) space, followed by partitioning it into bins. This mapping and binning process enables the dynamic splitting and merging of trajectories based on their progress along the identified CVs, thereby enhancing the sampling of rare events. The selection of appropriate CVs and binning schemes can significantly influence the algorithm’s performance. Voronoi polyhedra offer a convenient method to tessellate spaces of any dimensionality. This technique has been leveraged to enable adaptive binning in a hierarchical manner for the WE method.[15, 16] Based on the finite-temperature string method, Ref. 17 adaptively learns a string of Voronoi cells to restrict the WE sampling to a one-dimensional path embedded in a multi-dimensional space defined by some CVs. Later, a minimal adaptive binning (MAB) scheme[18] is proposed to automate the placement of bins along a chosen CV, which effectively surmounts large free energy barriers encountered during the WE simulation. Alternatively, recent advancements propose optimizing both bin construction and trajectory allocation based on a newly developed mathematical formulation.[19] Despite these strides towards more automatic binning schemes, the determination of CVs in WE simulations still heavily relies on expert intuition.
@@
Moreover, as an unbiased method, the WE method suffers from slow convergence. To speed up the convergence, Ref. 20 uses Gaussian accelerated molecular dynamics to generate an initial set of configurations. These configurations, weighted appropriately, are then utilized to seed the WE simulations, enhancing convergence. Similarly, Ref. 21 adopts a comparable approach, enhancing the initial state distribution for WE simulations by constructing a Markov state model (MSM) from short unbiased trajectories. Moreover, history-augmented Markov state models (haMSMs) have emerged as potent tools for estimating stationary distributions and rate constants from transient, unconverged WE data.[22] These estimates can then also serve as effective starting points for initiating new WE simulations, thereby accelerating the convergence of the simulation.
@@
Figure 1: Comparison of SPIB-based CVs and expert-based CVs. Expert-based CVs excel in extrapolation but may struggle with lifting intermediate state degeneracy. In contrast, SPIB-based CVs effectively distinguish distinct states and capture complex dynamics.
Figure 2: Hybrid approach flowchart and illustration. (a) Algorithm flowchart illustrating how the hybrid approach assigns replicas to bins. (b) Illustration of the hybrid approach combining expert-based CVs and data-driven SPIB-learned CVs for two instances of a protein conformation (top and bottom). Replicas are first mapped both to the expert-based and SPIB-learned CVs, denoted with subscripts $e$ and $s$ respectively. The top (bottom) replica is mapped to bin $i_{e}$ ($i^{\prime}_{e}$) based on expert-based bins and bin $i_{s}$ ($i^{\prime}_{s}$) based on SPIB-learned bins, as indicated by the arrows. Undersampled regions, defined by expert-based bins, are shaded in the bottom right figure. The total number of replicas mapped to these bins is represented by the number of circles. Since bin $i_{e}$ is not in an undersampled region, the top structure is assigned to bin $i_{s}$ (solid magenta arrow), regardless of the number of replicas in bins $i_{e}$ and $i_{s}$. However, since bin $i^{\prime}_{e}$ is in an undersampled region and contains fewer replicas (with only one replica) compared to bin $i^{\prime}_{s}$ (which has two replicas), the bottom replica is assigned to $i^{\prime}_{e}$ (solid green arrow).
@@ II.1 Weighted Ensemble (WE) Algorithm
In a typical MD simulation, the system spends most of its time in high-probability basins, and rarely leaves the basins to visit low-probability regions such as transition and metastable states. The WE algorithm [8, 9] adopts a splitting strategy using an ensemble of system copies (or “replicas”), each of which carries a statistical weight. These replicas are split or merged at a periodic time interval, called the resampling time $\tau$, guaranteeing an even coverage of the configuration space. Through this resampling procedure, the WE method is capable of enhancing the sampling of the low-probability regions. As these replicas are assigned statistical weights in a rigorous manner, in principle no bias is introduced in the dynamics. Thus, the WE algorithm can yield unbiased estimates of both the thermodynamic and kinetic properties of the molecular system.
@@ II.2 State Predictive Information Bottleneck (SPIB)
State predictive information bottleneck (SPIB)[23, 24] uses an information bottleneck-based protocol to learn CVs of desired dimensionality from time resolved trajectory data, while simultaneously identifying important metastable states. Previous studies have demonstrated its efficacy in analyzing complex molecular systems by constructing multi-resolution MSMs.[25, 24] Furthermore, SPIB has proven valuable in learning low-dimensional CVs for enhanced sampling methods. This approach accelerates diverse processes, including ligand permeation[26] and dissociation,[27, 28] protein conformational changes,[29, 30, 31] and crystal polymorph nucleation.[32, 33]
@@
SPIB adopts a structure akin to a variational autoencoder (VAE),[34] consisting of two neural networks: An encoder maps the input features, such as those characterizing a protein’s current configuration, to a low-dimensional CV space, while a decoder tries to predict some target from this space. However, unlike VAE, where the aim is to reconstruct the entire input, SPIB only requires the learned CVs to predict coarse-grained future state labels, namely which state the system will occupy after a pre-specified lag time $\Delta t$. The main advantage of such a simplification of the prediction task is that only the motion related to the transitions between different states will be captured by the learned CVs, while the fluctuations inside any metastable state will be ignored. Typically, the number and location of such states are not available a priori, but the SPIB method makes it possible to robustly estimate these on-the-fly.[23, 24] In this way, SPIB can autonomously cluster similar MD conformations into metastable states and accurately identify transition pathways between these states. This process allows SPIB to learn a low-dimensional CV space, where distinct states are separated into basins and connected by key transition pathways. Such clusterability of SPIB empowers the learned CVs to prioritize the representation of state-to-state transitions.
@@
By introducing a lag time $\Delta t$, SPIB acts as a “fast mode filter”, selectively identifying metastable states with significant Boltzmann weights and lifetimes surpassing $\Delta t$, while excluding modes equilibrating faster.[24] This enables SPIB to automatically pinpoint important metastable states in the system. $\Delta t$ serves as the minimum time resolution of interest. Aligning $\Delta t$ with the WE resampling time $\tau$—such as selecting $\Delta t$ as half of $\tau$—allows SPIB to filter out fast processes while retaining slower ones that cannot be sampled within $\tau$, making SPIB ideal for integration into WE simulations.
@@
The CVs learned by SPIB can be interpreted as a low-dimensional continuous embedding of MD conformations that preserves maximal information about state-to-state transitions, effectively capturing all transitions between identified states.[24] This is unlike commonly used dynamics-based dimension reduction methods such as tICA,[35, 36] which focus on learning a given number of slow processes. For instance, a 2D tICA captures only the two slowest processes while a 2D SPIB can accurately capture not just the two slowest, but all slow transition processes between the identified metastable states. This inherent distinction renders SPIB suitable for learning a low-dimensional CV space for systems featuring multiple metastable states of interest.
@@ II.3 SPIB Augmented Enhanced Sampling
A solution to this conundrum could involve an iterative approach that alternates between rounds of enhanced sampling and SPIB to gradually improve sampling.[37, 7] In the Reweighted Autoencoded Variational Bayes (RAVE) method,[38, 39, 26] the process starts with initial trajectories, and SPIB learns CVs, which may initially be suboptimal, to guide the subsequent enhanced sampling simulation to explore an increased amount of configuration space. Once a better sampling is collected, better CVs can then be learned. This looping of enhanced sampling and SPIB can continue until a satisfactory level of sampling is achieved.
However, as a deep neural network-based method, SPIB usually fails in extrapolation. This means that solely relying on SPIB to guide sampling may lead the system to get trapped in some metastable states, hindering effective exploration of new configurations. One solution to strengthen the extrapolability of such machine learned CVs is to use a linear encoder paired with a non-linear decoder in SPIB, as we have done in many previous applications.[27, 28, 29, 30, 31, 32, 33]
Here, we aim to explore an alternative approach by directly combining SPIB learned CVs with expert intuition based CVs (as illustrated in Fig. 1). This is primarily motivated by recognizing that while experts may not grasp all the intricate dynamics of complex molecular systems, they possess intuition about crucial, more global demarcators of state-to-state transitions. Hence, expert-based CVs like root-mean-square deviation of atomic positions (RMSD), radius of gyration (Rg), or specific distances and dihedral angles, remain prevalent in many current enhanced sampling studies, despite the emergence of many ML-based CV identification approaches. These expert-based CVs excel in discerning states of interest, exhibit good extrapolation, and provide valuable insights into system dynamics. However, they may not capture all transition details accurately and can suffer from degeneracy, thus hindering the identification of important intermediate states and exact transition pathways for complex systems. Moreover, due to the prohibitive cost associated with sampling high-dimensional spaces, these expert-based CVs also are constrained by their dimensionality, posing additional challenges in proposing effective CVs.
In contrast, ML-based CVs, such as those learned by SPIB, are inherently low-dimensional yet excel in discerning distinct conformations. They are adept at capturing dynamic details and identifying transition pathways, thereby significantly reducing the issue of degeneracy. However, despite their versatility and broad applicability, these methods often struggle with extrapolation. As a result, there is often a desire to incorporate system-specific intuition to guide exploration and CV identification processes in practical applications.
@@
Figure 5: Average free energy surfaces of alanine dipeptide from three separate 10 $\mu s$ runs of (b) the WE solely based on expert-based CVs and (c) the hybrid SPIB-WE, compared with (a) the reference 4 $\mu s$ brute-force simulations.
@@ II.4 A Hybrid Approach for CV construction for Weighted Ensemble
In this work, we adopt a uniform binning scheme for both expert-based CVs and SPIB-learned CVs. Therefore, combining expert-based and SPIB-learned CVs involves integrating bins derived from both sets of CVs. A simple strategy is proposed in Fig. 2. For instance, if a replica is assigned to bin $i_{e}$ based on the expert-based bins and to bin $i_{s}$ based on the SPIB-learned bins (where $i$ represents the bin index, and subscripts $e$ and $s$ denote expert-based and SPIB-learned, respectively), we need to determine when to use the expert-based binning and assign the replica to bin $i_{e}$ and when to use the SPIB-learned binning and assign the replica to bin $i_{s}$. First, we define the undersampled regions based on expert-based bins. Utilizing the trajectory data employed for SPIB training, we count the number of samples within each bin by mapping all samples to the expert-based bins, without considering sample weights. Subsequently, expert-based bins with counts below the average are identified as undersampled regions. This determination is updated after each SPIB training. Then, we assess whether the replica is in undersampled regions. If a replica is in undersampled regions and there are fewer replicas mapped to bin $i_{e}$ than to bin $i_{s}$, the replica is assigned to bin $i_{e}$; otherwise, it is assigned to bin $i_{s}$. This hybrid strategy effectively combines the strengths of both SPIB-based CVs and expert-based CVs. In well-sampled regions, SPIB-learned CVs effectively separate metastable states and highlight key transition pathways, offering reliable guidance for enhanced sampling and facilitating slow state-to-state transitions. Conversely, in undersampled regions, the hybrid approach utilizes both SPIB-based bins and expert-based bins to improve sampling of undersampled transition states and promote exploration of new regions. This approach facilitates effective navigation of complex systems, ensuring efficient and robust sampling across the configuration space of interest.
@@
The full workflow is summarized in Fig. 3. Initially, utilizing expert-based CVs and their corresponding binning scheme, $N_{0}$ iterations of WE simulations are conducted to gather initial trajectories. Next, we train the SPIB model on the gathered MD data, clustering the sampled MD conformations into metastable states. This process allows SPIB to effectively map explored regions to a low-dimensional CV space, where distinct states are separated into basins connected by key transition pathways. A more detailed discussion of the neural network architecture and model training is provided in the Supporting Information (SI). In this work, we employ a rectilinear grid with equal bin sizes to uniformly bin the SPIB-learned CV space. We implement a straightforward algorithm to automatically adjust the bin size, aiming to maintain a roughly constant desired number of occupied SPIB bins (see SI for details). These SPIB-based bins, in combination with the expert-based bins, are then used to guide further WE simulations. $N_{\text{update}}$ iterations of WE simulations are conducted using these hybrid bins. Subsequently, based on the newly collected data, a new SPIB model can be obtained. This iterative process between WE simulations and SPIB training continues until a satisfactory level of sampling is achieved, allowing for the final estimation of quantities of interest, such as rates or free energy differences.